# Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight

### v1

@@ 3 The Human Oversight Attack Surface
With a more wide-spread implementation, human oversight will also become an increasingly attractive target for malicious actors seeking to corrupt or disrupt AI-based operations. Such actors could be cybercriminals targeting high-risk AI systems for ransom or foreign governments attacking a country’s critical infrastructure in hybrid or open conflict. Furthermore, sufficiently powerful AI agents that are successfully attacked could themselves become malicious actors by attempting to evade oversight [69, 70]. Considering these threats, designing human oversight requires not only ensuring effective oversight [1] but also ensuring secure oversight.
Designing for secure human oversight demands an understanding of its attack surface. The following list of attack vectors inspired by the broader cybersecurity literature is not exhaustive but illustrates the breadth of the human oversight attack surface. It outlines attacks that target the four core requirements of effective human oversight: epistemic access, self-control, causal power, and fitting intentions. Table 1 summarizes these attack vectors and their relation to the requirements for effective human oversight.
Table 1: Attack vectors that target the effectiveness of human oversight. For each attack vector, we indicate the specific requirement for effective human oversight that the attack seeks to undermine.
Attack vectors Affected requirements of effective human oversight
Poisoning attacks Epistemic access, self-control
Adversarial attacks Epistemic access
Explainability attacks Epistemic access
(Distributed) denial of service attacks Causal power
Man-in-the-middle attacks Epistemic access, causal power
Vulnerability exploitation Epistemic access, causal power
Malware Epistemic access, causal power
Social engineering Epistemic access, self-control
Coercion Self-control
Bribery Fitting intentions
Insider threats Epistemic access, causal power, fitting intentions
3.1 Poisoning Attacks
Poisoning attacks target the core functionality of AI models, namely the association of patterns in data with outcomes (e.g., real-valued numbers in regression, class probabilities in classification, or new text/image data in generative AI) [71]. These attacks exploit the fact that modern AI models are often trained on large volumes of (semi-)automatically collected data, for example through web crawlers. Malicious actors can insert triggers, or predefined patterns, into such datasets. When models are trained on compromised data, they may consistently exhibit malicious behaviors whenever a trigger is presented [72].
Poisoning attacks can compromise both the epistemic access and the self-control of oversight personnel. They undermine epistemic access because poisoned samples containing hidden triggers cause AI responses that deviate from normal outputs. A carefully crafted trigger, such as an imperceptible pixel pattern in images [73], can prevent oversight personnel from recognizing the behavioral shift. Even when oversight personnel detect that something is amiss, hidden triggers may still make it difficult to understand the malicious behavior.
Poisoning attacks can also erode self-control. If triggered outputs cause the oversight interface to generate excessive micro-notifications or confirmation requests, malicious actors can exploit this to induce fatigue by overwhelming oversight personnel with trivial but time- and energy-consuming tasks. This not only reduces oversight efficiency by consuming resources but also increases the likelihood of errors once fatigue reaches a critical level, thereby undermining the security of human–AI interaction.
3.2 Adversarial Attacks
Adversarial attacks aim to induce misclassifications in the outputs of an AI model [71]. By feeding the system with adversarial examples, attackers seek either to fool the model or to reduce its accuracy and performance. These adversarial examples typically resemble normal inputs but include subtle perturbations that may not be perceptible to oversight personnel, leading to misclassifications aligned with the attacker’s goals [72].
Adversarial attacks primarily compromise the epistemic access of oversight personnel. A crucial part of understanding AI behavior is knowing how input patterns are mapped to outputs (e.g., which structures or features in an input image correspond to specific output classes). Without this knowledge, reliably determining whether the system behaves acceptably becomes exceedingly difficult. Adversarial perturbations obfuscate the relationship between inputs and outputs, reducing epistemic access to the model’s true behavior. As a result, oversight personnel may fail to detect unwanted model behavior or even system failures.
3.3 Explainability Attacks
Attacks on explainability specifically target explainability and interpretability approaches and tools. These attacks manipulate or deceive the evidence for a model’s reasoning. A key characteristic of such attacks is their ability to alter the explanation while keeping the model’s output unchanged. This means that the AI system may still produce the same output, but the explanation for that output is inaccurate or misleading [74].
Explainability attacks primarily undermine the epistemic access of oversight personnel. Misleading or inaccurate explanations prevent personnel from understanding the AI system’s true behavior. Attackers can exploit these attacks to fabricate behavior that serves their goals. Even with extensive querying, oversight personnel may be unable to detect such manipulations [75]. Moreover, a mismatch between outputs and explanations adversely affects personnel who rely on explanations to identify problematic AI behavior [25, 76].
These attacks can also induce a state of deceptive epistemic access, where the system appears transparent but provides inaccurate explanations. This scenario can be more dangerous than a “black box” system because it fosters unjustified trust and makes detecting incorrect outputs more difficult. Consequently, a key focus of explainability for human oversight must be ensuring the fidelity and robustness of explanations [77, 78].
3.4 (Distributed) Denial of Service Attacks
Denial of service (DoS) attacks refer to methods that render a system inaccessible [79]. The core principle of DoS attacks is to overwhelm a target system with more requests than it can handle, causing the system to become nonfunctional. Various techniques exist to perform a DoS attack, such as exploiting the TCP/IP protocol (SYN flooding), using the ping functionality (ping flooding), or targeting mail servers (mail flooding). A variation of DoS attacks is the distributed denial of service (DDoS) attack, which distributes requests across multiple systems rather than originating from a single source [80].
(Distributed) DoS attacks can undermine the causal power of oversight personnel. Modern computing systems are increasingly implemented as network-based, distributed systems—a standard likely to apply in human oversight settings, where the AI system runs on a cloud-based or on-premise server while oversight personnel interact with it via network protocols. (D)DoS attacks can make either the AI system or the human oversight interface inaccessible. This threat is especially severe for cloud-based AI systems: because the AI system and oversight personnel are physically separated, malicious actors can effectively disable oversight through (D)DoS attacks while the AI system continues to operate in an unsafe state or produces inaccurate outputs.
3.5 Man-in-the-Middle Attacks
An attack vector that can be used in computer networks is the Man-in-the-Middle (MitM) attack [44]. When successful, this strategy allows malicious actors to intercept and control data traffic between two or more systems, granting them access to the exchanged information and enabling them to manipulate it. Additionally, attackers can conceal their presence, tricking the involved systems into believing they are communicating securely with each other [44]. Various methods exist to carry out a MitM attack: ARP spoofing can be used when the attacker and target are on the same local area network (LAN), snarfing can be applied in WLAN-based attacks, and domain name system (DNS) cache poisoning is another possible approach.
MitM attacks can compromise the epistemic access and causal power of oversight personnel, depending on which direction the manipulated information travels. If the attacker intercepts and alters information from the AI system to the oversight personnel, epistemic access is undermined: the true behavior, functionality, or outputs of the AI system may remain hidden.
If the manipulated information flows from the human to the AI system, causal power is compromised. For example, if an attacker modifies commands issued by oversight personnel, they lose the ability to interact effectively with the AI system and cannot intervene in AI processes. This situation is particularly dangerous because the attacker can send manipulated responses back to the personnel, creating an illusion of control and misleading them into believing they are still effectively supervising the AI system.
3.6 Vulnerability Exploitation
Vulnerability exploitation refers to utilizing existing flaws in hardware and software for attacks, such as remote code execution or exfiltration of sensitive data. Since modern digital systems are complex combinations of hardware and software components and their respective supply chains, flaws can accumulate quickly and pose a significant threat. One of the most dangerous vulnerability exploitation is the zero-day exploit [81]. This term refers to the exploitation of a specific vulnerability within a hardware or software component that is unknown even to the manufacturer of the component. This gives the attacker the time advantage of day zero, before the manufacturer can detect, identify and fix the vulnerability.
Attacks based on vulnerability exploitation can target epistemic access or causal power. Attackers can use the vulnerabilities to get unauthorized access to the system and in this way manipulate the system output which the human oversight person uses to detect deviations in system behavior. If the attacker can obfuscate their actions from the human oversight personnel, the personnel are prevented from understanding the real functioning of the AI system. Similar effects can be achieved by manipulating other information that human oversight personnel uses to assess the AI system.
Causal power is affected by vulnerability exploitation if attackers prevent human oversight personnel from appropriately communicating with the AI system. This can be achieved by dropping commands send by the human oversight personnel or disabling crucial components of the human oversight software. Depending on the exploitation, attackers might be able to disguise the attack. For example, and similar to the example for MitM attacks, they could drop commands by human oversight personnel but fake a response by the AI system that suggests the command was successfully executed.
3.7 Malware
Malware is software that executes harmful functionality on an IT system, often without the user’s knowledge. Its potential targets and effects are highly diverse. Ransomware encrypts the file system and typically demands a ransom for decryption. Spyware extracts information from the infected system. These are just two of the most common types, but malware encompasses a wide range of threats designed for various purposes [82, 45].
Due to the diversity of malware functionality, it can compromise epistemic access and causal power of oversight personnel. Regarding epistemic access, malware can alter the core functionality of an AI system. For example, it could insert system prompts into deployed large language models (LLMs), significantly changing their behavior [83]. Malware can also intercept or manipulate data traffic from the AI system to oversight personnel, or modify, encrypt, or obfuscate system logs. These actions prevent personnel from understanding the system’s true functionality and therefore undermine epistemic access.
Causal power can be compromised if malware intercepts or manipulates data traveling from oversight personnel to the AI system. For instance, if malware prevents personnel from interacting with the AI system or locks them out of controlling its core functions, they lose the ability to intervene in system processes or outputs. Similarly, if malware secretly alters commands issued by oversight personnel, it further undermines their causal power.
3.8 Social Engineering
Social engineering refers to attacks that exploit human characteristics [84]. These attacks leverage traits such as fear, respect for authority, or the willingness to help, tricking victims into revealing personal information, disabling security measures, or installing malware. The widespread use of digital communication channels has provided malicious actors with effective and scalable tools for conducting social engineering attacks [85].
Social engineering can compromise epistemic access and self-control of oversight personnel. If an attacker impersonates individuals with technical expertise or authority (e.g., trusted colleagues, management, external experts) they can distort personnel’s understanding of relevant information needed to oversee AI systems. For example, an attacker might pretend to be a technician from the AI provider, convincing human oversight personnel that certain graphical user interface notifications indicate a bug. This could lead personnel to ignore important alerts or install malware disguised as a hotfix. Similarly, attackers impersonating management may convince oversight personnel that specific system notifications are non-critical, framing them as part of corporate processes. In both cases, social engineering feeds false information about the AI system’s functionality to oversight personnel, thereby undermining epistemic access.
Social engineering can also erode self-control. By impersonating senior colleagues or management, attackers can exploit perceived authority to pressure oversight personnel into taking specific actions. If the attack is credible, personnel might follow instructions against their better judgment, believing the directives come from a legitimate authority.
3.9 Coercion
Coercion is a broad concept that extends beyond IT security and involves using force, often through threats or violations, to influence a party. In the context of human oversight, coercion refers to pressure or force that alters the intended actions of oversight personnel.
Coercion undermines the self-control of oversight personnel. By influencing oversight personnel to act according to the instructions of a malicious actor rather than their own judgment, coercion renders human oversight ineffective. Even when oversight personnel detect a need for intervention, they may be unable to act according to their assessment. In extreme cases, a malicious actor could threaten the oversight personnel to gain control over their actions, effectively nullifying the intended benefits of human oversight.
3.10 Bribery
Bribery refers to the act of persuading a person, organization, or authority to act in accordance with the goals of a malicious actor, typically by offering monetary incentives, status, or similar benefits.
Bribery undermines the fitting intentions of human oversight personnel. Usually, the intentions of human oversight personnel should align with fulfilling their oversight duties, such as detecting inaccurate outputs or system failures [1]. A successful bribe can misalign these intentions, causing personnel to act in accordance with the malicious actor’s goals instead. Such goals might include damaging critical infrastructure or compromising the functioning of AI in high-risk applications, such as in hospitals.
3.11 Insider Threats
Insider threats encompass attack scenarios in which malicious actors infiltrate the organization they intend to compromise [86]. Such attacks share the common characteristic that the attacker either has direct access themselves or exploits an employee’s access rights, whether voluntarily or accidentally. Two pathways to insider threats are discussed under the sections on coercion (see 3.9) and bribery (see 3.10). Another common pathway is the exploitation of disgruntled employees.
Insider threats can undermine epistemic access, causal power, and fitting intentions of oversight personnel. They can erode epistemic access by altering or falsifying monitoring data from the AI system. Since human oversight personnel rely on this data to accurately assess AI behavior and intervene when necessary, this prevents them from understanding the system. Insider threats can also obfuscate AI system documentation by deleting or modifying key information, further impeding epistemic access.
Insider threats can also compromise the causal power of oversight personnel. For example, possible sequences of actions that can be taken as a response to detecting problematic system behavior can be ambiguous and the effectiveness of different sequences of actions can differ vastly. Attackers may be able to change organizational oversight policies on how to deal with inaccurate outputs or system malfunctions. In the worst case, such adapted policies can prevent oversight personnel from acting appropriately after detecting problematic system behavior which negatively affects causal power.
Another example of how insider threats may undermine causal power is by achieving and thereby exploit under-staffing. Redirecting resource allocation away from human oversight or preventing human oversight personnel from doing their work (e.g., by force) can reduce the workforce available to human oversight. While this requires substantial access and intervention capabilities, it can prevent oversight personnel from exercising causal power effectively.
Insider threats can additionally undermine the fitting intentions of oversight personnel. Malicious actors with insider access can manipulate the organizational environment to align human oversight personnel’s behavior with their goals. One method is positive punishment, such as establishing a “culture of fear” that discourages human oversight personnel in engaging in behavior that is aligned with their oversight duties. Another method is positive reinforcement, where incentive structures reward behaviors deemed advantageous by the malicious actor.
4 Hardening Human Oversight
Table 2: Hardening strategies are designed to mitigate the risks of (cyber)attacks on human oversight by malicious actors. Each strategy is mapped to the specific attack vectors it protects against.
Hardening strategy Targeted attack vector
Network management (D)DoS attacks
Intrusion Detection System (D)DoS attacks, MitM attacks, malware, insider threats
Encryption MitM attacks
Transparency All attack vectors
Governance principles and audit frameworks Vulnerability Exploitation, insider threats
Red Teaming Poisoning, adversarial, and explainability attacks, vulnerability exploitation
Training of human oversight personnel Social engineering, coercion, bribery
There are various hardening strategies to minimize the risk of malicious actors exploiting attack vectors. The following list of hardening strategies draws on cybersecurity research and practice. Similar to the list of attack vectors, this list is not exhaustive but is intended to inspire approaches for designing secure human oversight. Table 2 provides an overview of mitigation strategies and indicates which attack vectors they can address in the context of human oversight.
4.1 Network Management
As discussed in section 3.4, (D)DoS attacks overwhelm a network by generating excessive traffic. Defending against these attacks requires a multi-step approach [87, 88]. The first step is to deploy monitoring systems that analyze network traffic and trigger alerts if a potential (D)DoS attack is detected. The second step involves methods to distinguish between benevolent traffic (typically generated by humans) and malicious traffic (often produced by bots). Finally, filtering techniques can mitigate (D)DoS attacks by restricting or limiting what users – humans or bots – can do within the network.
Hardening strategy: Establish robust network management procedures within the human oversight architecture.
4.2 Intrusion Detection System
Intrusion detection involves the active monitoring of computers and networks to identify attacks and misuse. An intrusion detection system (IDS) typically combines multiple components and methods to perform this monitoring. The core technical components of an IDS include network or host sensors (depending on the IDS type), a database, a management module, and an analysis module. Host and network sensors monitor the system and collect the necessary information needed for the detection of attacks. The database needs to ensure that the data collected by sensors is stored with a sufficient degree of detail and for a sufficient amount of time. The management module provides all functionalities needed for configuring and calibrating the IDS. The analysis module provides all functionalities necessary for analyzing stored data and performing incident detection. A well-designed and maintained IDS can significantly enhance the cybersecurity of IT systems [89, 59]. Intrusion detection systems may counter multiple attack vectors, including (D)DoS attacks, MitM attacks, malware, and insider threats.
[…]

### v2

@@ 3 Threat modeling of human oversight
With a more wide-spread implementation, human oversight will also become an increasingly viable target for malicious actors seeking to corrupt or disrupt AI-based operations. Such actors could be cybercriminals targeting high-risk AI systems for ransom or foreign governments attacking a country’s critical infrastructure in hybrid or open conflict. Furthermore, sufficiently powerful AI agents that are successfully attacked could themselves become malicious actors by attempting to evade oversight [7, 51]. Considering these threats, designing human oversight requires not only ensuring effective oversight [72] but also ensuring secure human oversight.
In cybersecurity, the standard approach to identify, quantify and address potential flaws and vulnerabilities within an IT system is to perform threat modeling. Organizations and IT-professionals can use threat modeling to examine the IT system at hand from the viewpoint of malicious actors using a structured process. Threat modeling consists of four steps.
First, the scope of the tested application has to be determined, usually by creating a data flow diagram (DFD). Furthermore, the first step identifies entry points into the application, exit points of the application, privilege or trust boundaries, and assets that need to be protected. The second step involves determining threats to the tested application. In this work, the STRIDE [37] model is used for determining threats. STRIDE—an acronym for Spoofing, Tampering, Repudiation, Information disclosure, Denial of service, Elevation of privilege—is a well-established threat model. During the third step of threat modeling, countermeasures and potential hardening strategies are determined for identified threats. The fourth step consists of assessing the results of the previous steps and ensuring that no important points have been missed. For this manuscript, step four will be implicitly addressed in the discussion.
This section details the decisions we took in performing threat modeling for an IT-application representation of human oversight, and how we incorporated its socio-technical context into the analysis. We provide a guideline for researchers and practitioners who want to perform threat modeling for their real-world implementations.
3.1 Step 1 - The scope of human oversight
Figure 1: Data flow diagram (DFD) of the abstracted human oversight application used for threat modeling. Arrows indicate the directed paths of data through the application. Each arrow is labeled with the type of data that is transported. Circles indicate processes, two concentric circles indicate multiple subprocesses, and rectangles indicate external entities. Privilege boundaries are drawn with red dashed lines. HO = Human Oversight.
As mentioned above, the first step of any threat modeling is to collect information about the subject of the threat modeling, and thereby determine the scope of the work. A data flow diagram (DFD) shows different paths through an application and indicates the locations of privilege boundaries (also called trust boundaries). These boundaries indicate that movement along the path leads to a change of privileges, i.e., users and processes beyond the boundary have more rights to perform actions within the system like write data or run software. We extend the meaning of privilege boundaries to also include the right for intervention, meaning that entities with higher privilege can intervene on entities with lower privilege. The shape of nodes in a DFD are associated with specific meanings. A rectangle indicates an external entity, like a user that accesses the application. A circle indicates a single process. Two concentric circles indicate multiple subprocesses. Data flows are indicated by arrows, and privilege boundaries are indicated by dashed lines. Traditionally, DFDs can indicate data storage nodes with two parallel lines but we did not need this node type for the human oversight DFD.
Furthermore, DFDs are hierarchical in nature. This means that several different levels of DFDs can be created for an application. Each DFD focuses on a different part of the application. This approach has the advantage that every isolated DFD remains relatively simple. We use this property to compensate for the fact that we are not looking at a concrete implementation but a more abstract concept. Designing a high-level DFD allows us to omit implementation details via the two types of process nodes. Real-world implementations of human oversight can than fill in the details of the processes with lower-level DFDs.
3.1.1 The human oversight DFD
Figure 1 presents our DFD of the human oversight application. We identified three external entities. The users node includes all entities that interact with the AI system, either humans or other IT systems. The superior node indicates structures within the organization that have greater authority than the human oversight system. That could be for example the management of the organization. The external advisory board node indicates an external entity that defines appropriate actions and goals of human oversight, similar to an ethical advisory boards. Next, there are two multiple processes nodes in the DFD. One node represents the AI system, the other node represents the human oversight IT system. The human oversight IT system consists of all software that the human oversight personnel need to perform their duties. By choosing multiple processes nodes for the AI system and the human oversight IT system, we can create a top-level DFD that is valid for numerous implementations of human oversight. In case of threat modeling a concrete human oversight application in the future, these multiple processes nodes simply have to be specified with additional lower-level DFDs. Finally, the last identified node represents the human oversight personnel. We model this component of the human oversight application as a process to indicate the active role of humans in human oversight. Furthermore, the crucial role of human oversight personnel to the application is shown by the high connectivity of this node.
We identified several data flows within our theoretical human oversight application. Users send inputs to the AI system, which responds by sending outputs back to the users. Inputs can be any data that is used to interact with the AI system. Similarly, outputs can be arbitrary data that represents the output of the AI system (e.g., decisions or generated data points). There is a data flow between users and HO personnel indicating the possibility of users to directly send feedback and reports to the HO personnel. Such user reports are an important oversight signal in many applications. Another data flow connects the AI system with the human oversight personnel. While the AI system send operational parameters to the human oversight personnel (this is often called monitoring in the HO literature), the oversight personnel performs interventions on the AI system [41]. These can include commands to shut down, corrections of model parameters, or other interventions. The next data flow is located between human oversight personnel and human oversight IT system. While the personnel sends the operational parameter from the AI system to the human oversight IT system, the IT system uses these parameters to provide decision support for the oversight task. These will be transferred back to the human oversight personnel. The operational parameters could be directly send from the AI system to the human oversight IT system. However, This would not create a new privilege boundary crossing and, thus, can be viewed as similar to our DFD. We decided on the current DFD design to emphasis the importance of the human oversight personnel. There is a data flow between the human oversight personnel and the superior node. Since the superior node represents an entity of higher authority, the data flow labeling is chosen to represent this hierarchy. The human oversight personnel sends reports to the entity with higher authority while instructions (e.g., work instructions) are flowing the other way. Finally, data in form of guidelines flow from the external advisory board to the human oversight personnel.
Five privilege boundaries can be identified in the DFD. First, there is a privilege boundary between users and the AI system. While users should have minimal rights that are necessary to use the application, the AI system needs enough rights to fulfill its task. Additionally, there is a privilege boundary between users and HO personnel with a similar reasoning. The next privilege boundary lies between the AI system and the human oversight personnel. Being able to intervene on the AI system requires the HO system to have higher privileges that the AI system. The third privilege boundary lies between the human oversight personnel and the superior node. Since the superior node represents management of an organization, this node has the higher privilege context. The fourth privilege boundary lies between the human oversight personnel and the external advisory board. An external advisory board has no privileges within the system. That means we have five different privilege levels: advisory privilege, user privilege, AI privilege, HO privilege, and superior privilege.
3.1.2 Entry points
Table 1: A list of all identified entry points into the human oversight application. Each entry point has a unique identifier, a name, and a description. Furthermore, the column privilege levels indicate which privilege level are required at each entry point.
ID Entry point Description Privilege levels
Entry.1 User login Login procedure for users to interact with the AI system. This could be a login page reachable from the internet using a browser. Other possibilities are APIs, specialized user applications, or hardware components. user, AI, HO, superior
Entry.2 HO personnel login Login procedure for human oversight personnel to get access to the human oversight system. Possible options for this entry point are browser-based login pages, specialized applications, or hardware components. HO, superior
Entry.3 Access of superiors Capabilities of entities with higher authority (e.g., management of organizations) to gain access to the human oversight system. This can include organizational structures and intervention capabilities. superior
The next part of determining the scope of human oversight is to identify entry points into the application. The identification of entry points plays a crucial role in understanding possible paths that are open to a malicious actor to attack an application. Usually, this means to determine technical entry points like HTTP ports, login pages or functions, or search entry pages. Since we are not looking at a concrete implementation of human oversight, our analysis focuses on potential points of interest that need to be closely examined for real-world applications of human oversight. For every entry point, we also specify which privilege levels are required to use the entry point. Table 1 shows a summary of the entry points. The three identified entry points are the component used by users to access and use the AI system, the component used by human oversight personnel to access the human oversight application, and the component that allows entities with higher authority to access the human oversight application.
3.1.3 Exit points
Table 2: A list of identified exit points of the human oversight application. Each exit point has a unique identifier, a name, and a description. The last column indicates the corresponding entry point to each exit point, if such an entry point exists.
ID Exit point Description Corr. entry point
Exit.1 AI Output Communication channel that is used by the AI system to deliver outputs (e.g., predictions, decision, or generated data points) to users. Entry.1
Exit.2 HO authentication Response functionality by the human oversight application to deliver the results of the human oversight personnel’s login procedure. Entry.2
Exit.3 HO system report Communication channel between human oversight system and entities with higher authorities. This channel is used to deliver activity report and other requested information to superiors like the management of an organization. Entry.3
Some attacks require an exit point to achieve their goal, i.e., a gateway that can be used to ex-filtrate data from the application. Two more commonly known attacks worth mentioning here are cross-site-scripting vulnerabilities and information disclosure vulnerabilities. In general, exit points that are not sufficiently secured can be exploited by attackers to collect sensitive information. From a technical point of view, exit points often have a corresponding entry point. All exit points are summarized in Table 2. We identified three exit points for the human oversight application: the AI output provided to users, the component allowing human oversight personnel to access the human oversight application, and the component used to provide work reports to entities with higher authority.
3.1.4 Assets
Table 3: List of identified physical assets of the human oversight application. Each asset has a unique identifier, a name, and a description.
ID Name Description
PA.1 HO personnel credentials The keys that human oversight personnel uses to access the human oversight system.
PA.2 User data Depending on the concrete application, the human oversight system might store sensitive information about users, if this information is needed to determine whether an intervention has to be performed.
To complete the first step of threat modeling human oversight, we identify assets that need to be protected. Assets can be both physical and abstract. Physical assets are for example sensitive information like medical data of patients. Abstract assets can be understood as important properties and states, like the reputation of an organization. The lists of physical and abstract assets presented in this work are not exhaustive. Real-world applications of human oversight are likely to have additional assets.
Table 4: List of identified abstract assets of the human oversight application. Each asset has a unique identifier, a name, and a description.
ID Name Description
AA.1 Confidentiality Ensuring that sensitive information about the human oversight application, like interventions or decision protocols, cannot be accessed by unauthorized individuals.
AA.2 Integrity Ensuring the correctness of data and functionality of the human oversight application.
AA.3 Availability Ensuring that the human oversight application is functional at all times.
AA.4 Epistemic access Ensuring that human oversight personnel can sufficiently understand how the AI system works.
AA.5 Causal power Ensuring that the human oversight system can sufficiently intervene into the AI system.
AA.6 Self-control Ensuring that the human oversight personnel has sufficient control over their own actions.
AA.7 Fitting intentions Ensuring that the internal goals of human oversight personnel is aligned with the goals of the human oversight application.
We identify two physical assets (Table 3). The first are the credentials of human oversight personnel that are used to log into the human oversight system. These would allow an attacker to gain access with high privileges. Another physical assets are sensitive user information. Depending on the AI system, additional information about users might be necessary to ensure that the human oversight system has sufficient knowledge about a situation to decide whether or not to intervene. Envision an AI system that acts as a medical decision support system. Judging the validity of decisions made by such a system requires detailed knowledge of patient-specific information. In general, such sensitive information has to be protected against unauthorized access.
We identified seven abstract assets (Table 4). The first three assets that need to be protected are defined by the basic focus of cybersecurity, the CIA triad. This concept describes three main aspects of IT systems that have to be protected. The first aspect is the confidentiality of information within the system (denoted by the letter ’C’). Confidentiality is closely connected with the physical assets identified above but also covers appropriate management of read access within the application. The second aspect is the integrity of data and functionality of the IT system (denoted by the letter ’I’). For human oversight this can include for example ensuring correctness of AI operational parameters or correct functioning of software components. The last aspect covered by the CIA triad is availability (denoted by the letter ’A’). For example, it needs to be ensure that the human oversight personnel can access the AI system or the human oversight IT system. Furthermore, the obligations associated with human oversight can only be fulfilled, if such oversight can be considered effective. Therefore, the four requirements of effective oversight [72] have to be added to the list of abstract assets. The first asset is epistemic access, which means that human oversight personnel has to be able to understand the functioning of the AI system. The next asset is causal power, which means that the human oversight system has to be able to perform interventions on the AI system. The third requirement is self-control, which means that the human oversight personnel of our application can make decisions without malicious outside influence. Finally, the fourth requirement is fitting intentions. Mapping this requirement onto our human oversight application means that the goals of the human oversight personnel have to be aligned with the overall goals of human oversight.
Based on the result of the first step of threat modeling, we now continue to identify security threats that need to be considered when implementing a human oversight application.
3.2 Step 2 - Security threats
We use the established threat model STRIDE (Spoofing, Tampering, Repudiation, Information disclosure, Denial of services, Elevation of privilege) to identify security threats to the human oversight application. STRIDE allows us to examine the application from an attacker’s perspective and to identify vulnerabilities with the attacker’s goals in mind.
3.2.1 Spoofing
The term spoofing summarizes threat actions that are aimed at accessing and using credentials without permission. The goal is to interact with a system while assuming a false identity. This can allow attackers to gain unauthorized access to sensitive data or functionalities.
The obvious target for spoofing attacks in our human oversight application is the human oversight personnel and, thereby, using entry point HO personnel login (Entry.2) for exploitation. Human oversight personnel has to have far reaching privileges within the human oversight application, thus, gaining access through the HO personnel login can be a high priority target for attackers. The first set of tools readily available to attackers to pursue this target are the different methods from social engineering. This include water holing, baiting, phishing, vishing, smishing, and others. Any of the mentioned approaches can be used to steal credentials and use them for unauthorized access to the human oversight application with considerable privileges. If there is insufficiently secured remote access to the human oversight application, attackers can perform a man-in-the-middle (MITM) attack to get transferred information, which in this case are credentials. The term MITM describes attacks that rely on the attacker’s ability to secretly intercept, relay, and possibly alter communication between two parties. The abstract assets that are vulnerable to social engineering attacks are confidentiality (AA.1) and self-control (AA.6). The vulnerability of confidentiality is self-evident. Since the credentials themselves are considered confidential information, attacker access to credentials is a breach of confidentiality. The vulnerability of self-control becomes clear when looking at the goal of spoofing attacks: accessing an application while assuming a false identity. Attackers that use stolen credentials and are able to use these credentials to log into the human oversight application can act as human oversight personnel. For the system, all malicious actions taken will look like actions by human oversight personnel. Therefore, the human oversight personnel is no longer in control of their own actions. The physical asset HO personnel credentials (PA.1) is a target of these types of attacks.
Another possible direction of spoofing attacks has emerged with the rise of more and more powerful agentic AI. This direction comes from the AI agents themselves. Research has suggested that current state-of-the-art large language models (LLMs) have the ability to do scheming [51]. This term summarizes different misaligned behavioral patterns shown by LLMs that are meant to pursue objectives that differ from expectations of human users. With insufficiently secured model context protocol (MCP) servers, AI agents can try and use available tools to access credentials of human oversight personnel. Using these credentials, the AI agent can act within the human oversight system acting as the human oversight personnel. In this scenario, human oversight would be completely bypassed by the AI agent. Obviously, one abstract asset that is attacked in this scenario is integrity (AA.2). An agentic AI bypassing human oversight renders the functionality of the human oversight application useless. Furthermore, self-control (AA.6) is also attacked since a successful attack results in another entity being in control of the human oversight personnel’s actions. One might argue that also the assets confidentiality (AA.1) and HO personnel credentials (PA.1) are attacked. However, we would argue that misaligned behavior of AI agents do not expose confident data to the outside of the application. Therefore, these assets are not attacked.
3.2.2 Tampering
The term tampering summarizes attacks that are aimed at maliciously modifying a system. This can include changing persistent data, altering the system’s behavior, or messing with communication in transit.
One target for tampering attacks is the communication from users to the AI system. Malicious actors can use (il)legitimately obtained user credentials to send (indirect) prompt injections to the AI system. Additionally, MITM attacks can be used to modify legitimate traffic coming from users with (indirect) prompt injections. Such attacks can cause AI systems to produce unexpected behavioral patterns. First asset vulnerable to this attack is integrity (AA.2) since (indirect) prompt injection cause unexpected behavior of AI systems and, thus, attack the functional integrity of the system. The second vulnerable asset is epistemic access (AA.4). Since tampering attacks alter the behavior of AI systems and can go undetected, the knowledge of the HO system on how the AI system works gets eroded. If the AI system has the rights to write data, the physical asset user data (PA.2) can also be vulnerable to these attacks. Malicious actors can try to exploit these rights and trigger the AI system to change stored user data.
Another possible tampering attack is targeting the supply chain to poison training data or directly modify model parameters. Poisoning attacks can embed back doors into AI models, which can be exploited once the model is deployed. Again, backdooring AI models and tampering with parameters erode the HO system’s fundamental understanding of how the AI system functions. Thus, epistemic access (AA.4) is a target of these attacks. Furthermore, integrity (AA.2) is also a target, since the attacks aim to compromise the functional integrity of the AI system.
3.2.3 Repudiation
The term repudiation summarizes attacks that are aimed at performing prohibited operations within an IT system. These can exploit lacking capabilities to track operations within a system.
The first attack vector that can be exploited for repudiation are adversarial attacks and (indirect) prompt injections. Using the communication channel between AI system and users, malicious actors can try to circumvent guardrails embedded into AI systems to trigger forbidden operations. In the worst case, sensitive data can be extracted exploiting the lethal trifecta [77] together with exit point AI output (Exit.1). If the AI system has access to log files, malicious actors can try to hide these forbidden operations. Therefore, (indirect) prompt injections can be used to attack both physical assets, HO personnel credentials (PA.1) and User data (PA.2). Furthermore, the abstract assets confidentiality (AA.1), integrity (AA.2), and epistemic access (AA.4) are vulnerable to these attacks.
LLM’s ability to perform scheming can also open the human oversight application to repudiation attacks. Scheming can convert an AI system into an insider threat by triggering forbidden operations due to misaligned behavior. Again, these attacks can be obscured, if the AI system has the ability to alter log files. Epistemic access (AA.4) and integrity(AA.2) are vulnerable to this form of attack, since scheming prevents human oversight personnel to fully understand the functioning of the AI system and the functionality of the AI system is compromised. Depending on how extensive the misaligned AI system can operate within the application, the asset causal power can also be vulnerable. An AI system with sufficient rights can potentially limit the intervention capabilities of the human oversight personnel.
Another set of attack vectors for repudiation attacks are directly targeted at the human oversight personnel. Bribery and coercion can be used by malicious actors to force human oversight personnel to perform prohibited operations. Again, being able to alter the log files allows human oversight personnel to hide these forbidden operations. The abstract assets confidentiality (AA.1) and integrity (AA.2) are vulnerable to these types of attacks, since both bribery and coercion can be used to extract sensitive information from the system and maliciously alter persistent data and functionalities. Furthermore, coercion attacks self-control (AA.6) due to the outside force that determines the actions for the human oversight personnel. Bribery attacks fitting intentions (AA.7), since the original goal of the human oversight personnel gets overridden with monetary or otherwise self-serving interests.
3.2.4 Information disclosure
The term information disclosure summarizes attacks that aim at reading files without permission or reading data in transit. Since all attacks that fall into this category are explicitly targeting the disclosure of sensitive information, the asset that is mainly targeted by these attacks is confidentiality (AA.1). Furthermore, the two identified physical assets HO personnel credentials (PA.1) and user data (PA.2) are sensitive information and, thus, can also be targeted by information disclosure attacks.
The two data flows that will be the main target of information disclosure attacks are between AI and human oversight personnel, and human oversight personnel and IT system. Both data flows contain valuable information about the inner workings of the human oversight applications and, hence, might be interesting for attackers. The data flow between users and AI system could also be attacked. However, this data flow contains mainly information about the inner workings of the AI system. This data flow plays only a secondary role for the threat modeling of human oversight. A standard attack vector for data in transit are MITM attacks.
3.2.5 Denial of services
[…]